An R-C circuit is driven by an alternating voltage of amplitude 110 V and frequency ω. Define Vout to be the amplitude of the voltage across the capacitor. The resistance of the resistor is 1000 Ω, and the capacitance of the capacitor is 1.00 μF .

Part A What is Vout if ω=10radians/s? Express your answer numerically, in Volts.

V Part B What is Vout if ω=105radians/s? Express your answer numerically, in volts.
Answer #1

Given, Vo = 110V R=1000R (= IME = 1x10bf part - A - YAD/ d Xc = uc = 105 = Jox 1x106 Impedence of the circuit, Z J (1000) + (Part-B we 105 rad/s we = =9523.812 1058 10-6 z = J + xe 7 (1000) (9523.81? 9576.16 s the cisenit, Current through To Z 9576.1Thanks